please help.



2 to place and Say - for i and F to twe deimal places. Round your anowers for a and b to three demal places)
media%2F51a%2F51a94e14-0e51-49a6-8be6-af
Answer #1

a) Draw a scatter diagram displaying the data.

Answer:

60 ך 50 40 20 10 0 0 3 4 5 7

b)

\sum x=13

\sum y=151

\sum x^{2}=65

\sum y^{2}=6047

\sum xy=403

r=-0.988

c.

\bar{x}=3.25

\bar{y}=37.75

\hat{y}=50.286+(-3.857)x

35 50 4 5 4D 35 30 25 20 30 45 40 SD 25 20 13p 50 13 35 35 30 23 23

e)

r^{2}=0.976

explained=97.6\%

un-explained=2.4\%

(f)

\hat{y}=50.286+(-3.857) \times 3

=50.286-11.571

=38.72%